in a laboratory,1.55g of organic compound containing C, H, and O is combusted for analysis combustion resulted in 1.45g of carbon II oxide and 0.890g of steam. what is the empirical formular of the organic compound


1
Expert's answer
2022-02-08T21:29:02-0500

Mass of O2 = 2.55 - 1.45 - 0.890 = 0.21 g .

Moles of O2 = 0.21 / 32 = 0.006

Moles of CO2 = 0.032

Moles of H2O= 0.890 / 18 = 0.049444


Then find molar ratio of CO2 and H2O.

So, ratio is 2 x : y =

moles of CO2 / moles of H2O .

Where x and y are atomicity of atom of c and H.